# Precise calculation of the EFT likelihood with primordial non-Gaussianities

### v1

@@ 3 Generalizing the likelihoods to the non-Gaussian regime
One significant advantage of the path-integral approach to the LSS is that different cosmological effects can be systematically encoded as the corresponding effective terms of the partition functions, in accordance with the spirit of Effective Field Theory [27, 28, 33]. The objective of this section is to introduce PNG and elucidate how their effects can be incorporated into the partition functions. As we shall see, at leading order, the implementation of PNG is equivalent to introducing a cubic interaction term in the initial matter field within the partition functions. This term is proportional to both $f_{\rm NL}$ and $\delta_{\rm in}^{3}$, which requires careful treatment in the computation of the likelihoods. At the same time, this feature allows us to employ perturbative methods as in section 4 to compute the leading-order $f_{\rm NL}$ corrections to the likelihoods.
@@ 3.1 Introducing primordial non-Gaussianities
$\varphi_{\rm NG}(\boldsymbol{k})=\varphi(\boldsymbol{k})+f_{\rm NL}\int_{\boldsymbol{p}}K_{\rm NL}(\boldsymbol{p},\boldsymbol{k}-\boldsymbol{p})\left[\varphi(\boldsymbol{p})\varphi(\boldsymbol{k}-\boldsymbol{p})-P_{\rm G}(p)\delta_{D}(\boldsymbol{k})\right]+\cdots\,,$ (3.1)
where $\varphi$ is a Gaussian random variable, $P_{\rm G}(k)\equiv\left\langle\varphi(\boldsymbol{k})\varphi(-\boldsymbol{k})\right\rangle^{\prime}$, and the $\cdots$ represents higher-order non-Gaussian contributions. Throughout this paper, we will assume that the field without the subscript “NG” corresponds to a Gaussian field and will only consider the leading-order non-Gaussianities. $K_{\rm NL}$ represents the kernel function dependent on the inflationary models. This expression leads to a non-vanishing primordial bispectrum
@@
where “perms” denotes the cyclic permutations of the three momenta $k_{1}$, $k_{2}$, $k_{3}$. In the squeezed limit, which is of critical importance for the study of galaxy clustering, the non-singular kernels can be uniquely determined  [^5] It is worth noting that the existence of particles beyond the standard model with masses comparable to the Hubble scale during inflation can potentially induce non-analytic features in the squeezed limit through the cosmological collider mechanism [46, 8]. [13]. In this case, we have
@@
Under the assumption of spatial homogeneity and isotropy, the kernel $K_{\rm NL}(\boldsymbol{k}_{\ell},\boldsymbol{k}_{s})$ can only depend on the magnitude of the two momenta $k_{\ell}$, $k_{s}$, and their relative angle $\hat{\boldsymbol{k}}_{\ell}\cdot\hat{\boldsymbol{k}}_{s}$. Furthermore, due to the $\boldsymbol{k}_{s}\rightarrow-\boldsymbol{k}_{s}$ symmetry of the bispectrum that arises in the squeezed limit, the kernel function can be expanded in a series of even-order Legendre polynomials [26, 13]
@@ 3.2 The partition functions with primordial non-Gaussianities
In this expression, we have omitted the dependence of the transfer function $T$ on the conformal time $\tau$, which should be kept in mind. The normalization condition of this function is $T(k)\rightarrow 1$ for $k\ll k_{\rm eq}$. Then we employ the gravitational Poisson equation in Fourier space [36]
@@
where $H_{0}$ is the present-day Hubble constant, $\Omega_{m,0}$ is the fractional matter density and $D$ the growth factor. From these two expressions, we obtain the relation between the primordial potential and the initial matter field
@@
Analogously, we define the non-Gaussian matter field $\delta_{\rm in,NG}$, such that eq. (3.7) naturally induces its expansion into the following form
$\delta_{\rm in,NG}(\boldsymbol{k})=\delta_{\rm in}(\boldsymbol{k})+f_{\rm NL}\int_{\boldsymbol{p}_{1},\boldsymbol{p}_{2}}\delta_{D}(\boldsymbol{k}-\boldsymbol{p}_{12})K_{\rm NL}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})\frac{M(|\boldsymbol{p}_{1}+\boldsymbol{p}_{2}|)}{M(p_{1})M(p_{2})}\delta_{\rm in}(\boldsymbol{p}_{1})\delta_{\rm in}(\boldsymbol{p}_{2})\,.$ (3.8)
The contributions from higher-order PNG can be obtained through an analogous procedure. This expression allows for a straightforward incorporation of arbitrary order PNG into the partition functions via appropriate corrections to the initial matter field $\delta_{\rm in}(\boldsymbol{k})$. For example, when we want to generalize the following matter partition function as presented in [23]
@@
where $\mathcal{P}[\delta_{\rm in}]$ is the initial probability distribution which is usually assumed to be Gaussian, and $P_{\epsilon_{m}}(k)$ denotes the power spectrum of the noise in the matter density field to cancel the UV dependence of the loop diagrams [23]. At this time all that required is to substitute the non-Gaussian statistics of $\delta_{\rm in}$ eq. (3.8) into the probability distribution
@@
$-f_{\rm NL}\int_{\boldsymbol{p}_{1},\boldsymbol{p}_{2},\boldsymbol{p}_{3}}\left[\delta_{D}(\boldsymbol{p}_{123})K_{\rm NL}(\boldsymbol{p}_{2},\boldsymbol{p}_{3})\frac{M(p_{1})}{M(p_{2})M(p_{3})}\frac{\delta_{\rm in}(\boldsymbol{p}_{1})}{P_{\rm in}(p_{1})}\delta_{\rm in}(\boldsymbol{p}_{2})\delta_{\rm in}(\boldsymbol{p}_{3})\right]\,.$ (3.10)
@@
×$\times\exp\left\{-\frac{1}{3!}\,f_{\rm NL}\int_{\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2}}\delta_{D}(\boldsymbol{k}+\boldsymbol{p}_{12})B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\delta_{\rm in}(\boldsymbol{k})\delta_{\rm in}(\boldsymbol{p}_{1})\delta_{\rm in}(\boldsymbol{p}_{2})\right\}\mathcal{P}[\delta_{\rm in}]\,,$ (3.11)
@@
Note that the sign convention adopted here is consistent with [53], but opposite to that in [37]  [^6] In fact, both sign conventions are equally valid and yield the correct $n$-point statistics, as explicitly shown in eq. (4.2) of [53] and eq. (B.5) of [37]. The difference between them merely corresponds to a redefinition of the “coupling parameter” $B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\rightarrow-B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$, and after matching with the $n$-point correlation functions, this distinction does not affect any physical predictions.. In this context, the coupling parameter $B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$ does not exhibit symmetry under cyclic permutations of $k,p_{1},p_{2}$, but instead retains symmetry only under the exchange of $p_{1}$ and $p_{2}$. At the same time, $\mathcal{P[\delta_{\rm in}]}$ is still modeled as a Gaussian probability distribution, despite the fact that the underlying primordial statistics are no longer Gaussian. It is evident that, after incorporating PNG, the leading order correction in $f_{\rm NL}$ effectively manifests as an additional interaction term in the partition function. Moreover, as illustrated in [37], this term can be absorbed into the term proportional to $J$ after performing a shift of the integration variable, which is consistent with the arguments in quantum field theory. Following the same procedure, we can derive the expression for the galaxy partition function with PNG
@@
×$\times\exp\left\{\int_{\boldsymbol{k}}\left[\frac{1}{2}P_{\epsilon_{m}}(k)J(\boldsymbol{k})J(-\boldsymbol{k})+J_{g}(\boldsymbol{k})\delta_{\rm g,fwd}[\delta_{\rm in}](-\boldsymbol{k})+J(\boldsymbol{k})\delta_{\rm fwd}[\delta_{\rm in}](-\boldsymbol{k})+\cdots\right]\right\}$
×$\times\exp\left\{-\frac{1}{3!}\,f_{\rm NL}\int_{\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2}}\delta_{D}(\boldsymbol{k}+\boldsymbol{p}_{12})B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\delta_{\rm in}(\boldsymbol{k})\delta_{\rm in}(\boldsymbol{p}_{1})\delta_{\rm in}(\boldsymbol{p}_{2})\right\}\mathcal{P}[\delta_{\rm in}]\,.$ (3.13)
Compared to eq. (3.11), we have not only extended the current to separately correspond to the matter and galaxy fields, but have also generalized the stochastic term, then have (a) the stochasticity for galaxies $P_{\epsilon_{g}}\sim k^{0}$, (b) the cross stochasticity between galaxies and matter $P_{\epsilon_{g}\epsilon_{m}}\sim k^{2}$, and (c) the matter stochasticity $P_{\epsilon_{m}}\sim k^{4}$ (where $\sim$ represents their leading-order contributions). When evaluating the joint likelihood in the large-scale limit, the matter stochasticity can usually be safely omitted, since it involves higher-order momentum dependence, which allows us to take $P_{\epsilon_{m}}\rightarrow 0$ [23].
In summary, in both cases, the effects of PNG can equivalently be reformulated as an interaction term under the Gaussian initial conditions, thereby enabling the direct application of all analyses developed within the Gaussian framework. For example, in the current context, the matter likelihood can still be expressed as the functional Fourier transform of the matter partition function
@@
Here, we provide a brief comparison between the path-integral formulation and the standard perturbation theory (SPT) approach (see, for example, in [54]) in the context of incorporating PNG. In SPT, it is unfeasible to directly incorporate contributions from PNG in the evolution equations. Instead, one needs to compare the predicted bispectrum from the solutions with that in the non-Gaussian case, thereby extracting the additional expansion of the field, as illustrated in [53]. However, in the path-integral formulation, we can first postulate the effective terms induced by PNG, and then compare the predicted correlation functions with experimental data to constrain the magnitudes of the parameters in these effective terms. We argue that the path-integral approach offers a distinct advantage in addressing the issue of PNG, particularly when considering higher order non-Gaussianities. For example, in [37], the authors employed this approach to derive the renormalization group equations of LSS  [^7] In the context of EFTofLSS, the renormalization group equations have a meaning distinct from that in quantum field theory. They represent the running of physical quantities as the momentum cutoff is varied [29, 51, 52, 37], and can therefore also be interpreted as the evolution equations. in the presence of PNG, indicating that the incorporation of PNG into the evolution equations is feasible. A further example is provided by the comparison between the formulae in [55] and [53]: the former demonstrates that the probabilistic evolution equation coincides with the SPT evolution equation, whereas the latter establishes that the path-integral formulation offers a natural framework for incorporating PNG. At the same time, we can observe from the expressions for the two likelihoods eqs. (3.14) and (3.15) that, after the inclusion of PNG, both likelihoods differ from the Gaussian counterparts only by the addition of an effective term. This enables us to build upon the previous results in the calculation while incorporating the additional terms proportional to $f_{\rm NL}$. In the next section, we will employ the saddle-point expansion method to compute these two likelihoods and demonstrate how the presence of PNG gives rise to additional nontrivial contributions to the conditional likelihood.
@@ 3.3 Dimensional analysis
As we can see, the expressions for the two likelihoods eqs. (3.14) and (3.15) involve numerous terms to be handled. Evaluating these likelihoods directly with all terms included poses a considerable challenge. Therefore, in this subsection we perform a dimensional analysis to investigate the relative contributions of each term to the likelihoods at large scales. This can be achieved by performing a momentum rescaling analogous to the procedure outlined in [23], with $\boldsymbol{k}=b\boldsymbol{k}^{\prime}$. Then, in the limit $b\rightarrow 0^{+}$, the behavior of different terms under this rescaling will reveal their relevance (or irrelevance) to the large-scale universe. Under this rescaling, the integration measure ${\rm d}^{3}k$ changes to $b^{3}{\rm d}^{3}k^{\prime}$, and the power spectrum $P_{\rm in}$ scales as $P_{\rm in}(k)=b^{n_{\delta}}P_{\rm in}(k^{\prime})$, where $n_{\delta}$ is the order of momentum in the initial power spectrum. Then, the terms in the partition function that are quadratic in $X_{g}$ and $\delta_{\rm in}$ remain invariant if we redefine
@@
These are the results of the rescaling in momentum space, and the scaling behavior of these operators in real space can be obtained via the inverse Fourier transform
@@
To determine the rescaling of the non-Gaussian effective vertex, it is also necessary to study the scaling behavior of the non-Gaussian parameter $B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$ in eq. (3.12). We begin in momentum space, given the asymptotic behavior of the transfer function $T(k)\rightarrow 1$ in the limit $k\rightarrow 0$, it can be treated as dimensionless with respect to momentum at large scales, and accordingly exhibits a trivial scaling behavior, namely $b^{0}$. In this way, $M(k)$ can be approximated to have a scaling behavior of $k^{2}$ in the relevant regime. Meanwhile, for the lowest-order momentum-dependent PNG — corresponding to the spin-0 case discussed in [13] — the associated kernel function can also be either considered momentum dimensionless. After substituting the momentum dependence of the initial power spectrum, the rescaling behavior of this parameter in real space is given by
@@
Based on these expressions, we can assess the contributions of the different terms in the partition function eq. (3.11) at large scales. We present the result of the rescaling of these terms and the evolution of their scaling dimensions with respect to $n_{\delta}$ in Table 1 and Figure 1, respectively. It is evident that for $n_{\delta}=-1.7$, both stochastic terms $P_{\epsilon_{g}}JJ$ and $P_{\epsilon_{g}\epsilon_{m}}J_{g}J$ exhibit rescaling with a higher-order dependence on $b$ relative to PNG effective term, thereby yielding subdominant contributions at large scales. This is precisely why, in [37], the authors were able to disregard the stochastic terms and focus primarily on the non-Gaussian contributions when deriving the renormalization group equations incorporating PNG. At the same time, we can observe that in our case, the $P_{\epsilon_{g}\epsilon_{m}}J_{g}J$ term generally contributes the least to the partition function. This is because the leading-order momentum dependence of $P_{\epsilon_{g}\epsilon_{m}}(k)$ scales as $k^{2}$, resulting in a stronger suppression compared to other operators. Consequently, it is reasonable to neglect the higher-order contributions from $P_{\epsilon_{g}\epsilon_{m}}$ when calculating the likelihoods. The perturbative properties of $f_{\rm NL}$ and $P_{\epsilon_{g}\epsilon_{m}}$ will constitute the central framework for our detailed calculations in the next section, enabling us to extract the leading-order additional terms associated with the non-Gaussian contributions.
4 Precise calculation of the EFT likelihood

### v3

@@ 3 Generalizing the likelihoods to the non-Gaussian regime
One significant advantage of the path-integral approach to the LSS is that different cosmological effects can be systematically encoded as corresponding effective terms in the partition functions, in accordance with the spirit of Effective Field Theory [28, 29, 34]. The objective of this section is to introduce PNG and elucidate how their effects can be incorporated into the partition functions. As we shall see, at leading order, the implementation of PNG is equivalent to introducing a cubic interaction term in the initial matter field within the partition functions. This term is proportional to both $f_{\rm NL}$ and $\delta_{\rm in}^{3}$, which requires careful treatment in the computation of the likelihoods. At the same time, this feature allows us to employ perturbative methods, as in section 4, to compute the leading-order $f_{\rm NL}$ corrections to the likelihoods.
@@ 3.1 Introducing primordial non-Gaussianities
$\varphi_{\rm NG}(\boldsymbol{k})=\varphi(\boldsymbol{k})+f_{\rm NL}\int_{\boldsymbol{p}}K_{\rm NL}(\boldsymbol{p},\boldsymbol{k}-\boldsymbol{p})\left[\varphi(\boldsymbol{p})\varphi(\boldsymbol{k}-\boldsymbol{p})-P_{\rm G}(p)\hat{\delta}_{D}(\boldsymbol{k})\right]+\cdots\,$, (3.1)
where $\varphi$ is a Gaussian random variable, $P_{\rm G}(k)\equiv\left\langle\varphi(\boldsymbol{k})\varphi(-\boldsymbol{k})\right\rangle^{\prime}$, in which $\left$$\cdots\right$$^{\prime}$ indicates that an overall momentum-conserving delta function has been dropped. We have also defined $\hat{\delta}_{D}(\boldsymbol{k}_{1}+\boldsymbol{k}_{2})=(2\pi)^{3}\delta_{D}(\boldsymbol{k}_{1}+\boldsymbol{k}_{2})$. The $\cdots$ in eq. (3.1) represents higher-order non-Gaussian contributions. Throughout this paper, we assume that the fields without the subscript “NG” correspond to Gaussian fields and only consider the leading-order non-Gaussianities. $K_{\rm NL}$ represents the kernel function, which is dependent on the inflationary models. This expression leads to a non-vanishing primordial bispectrum
@@
where “perms” denotes the cyclic permutations of the three momenta $k_{1}$, $k_{2}$, and $k_{3}$. In the squeezed limit, which is of critical importance for the study of galaxy clustering, the non-singular kernels can be uniquely determined [13]. [^6] It is worth noting that the existence of particles beyond the standard model with masses comparable to the Hubble scale during inflation can potentially induce non-analytic features in the squeezed limit through the cosmological collider mechanism [53, 8]. In this case, we have
@@
Under the assumption of spatial homogeneity and isotropy, the kernel $K_{\rm NL}(\boldsymbol{k}_{\ell},\boldsymbol{k}_{s})$ can only depend on the magnitudes of the two momenta $k_{\ell}$, $k_{s}$, as well as their relative angle $\hat{\boldsymbol{k}}_{\ell}\cdot\hat{\boldsymbol{k}}_{s}$. Furthermore, due to the $\boldsymbol{k}_{s}\rightarrow-\boldsymbol{k}_{s}$ symmetry of the bispectrum that arises in the squeezed limit, the kernel function can be expanded in a series of even-order Legendre polynomials [44, 13]
@@ 3.2 The partition functions with primordial non-Gaussianities
In this expression, we have omitted the dependence of the transfer function $T$ on the conformal time $\tau$, which should be kept in mind. The normalization condition of this function is $T(k)\rightarrow 1$ for $k\ll k_{\rm eq}$. We then employ the gravitational Poisson equation in Fourier space [43]
@@
where $H_{0}$ is the present-day Hubble constant, $\Omega_{m,0}$ is the fractional matter density, and $D$ is the growth factor. From these two expressions, we obtain the relation between the primordial potential and the initial matter field
@@
Analogously, we define the non-Gaussian matter field $\delta_{\rm in,NG}$ such that eq. (3.7) naturally induces its expansion into the following form
$\delta_{\rm in,NG}(\boldsymbol{k})=\delta_{\rm in}(\boldsymbol{k})+f_{\rm NL}\int_{\boldsymbol{p}_{1},\boldsymbol{p}_{2}}\hat{\delta}_{D}(\boldsymbol{k}-\boldsymbol{p}_{12})K_{\rm NL}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})\frac{M(|\boldsymbol{p}_{1}+\boldsymbol{p}_{2}|)}{M(p_{1})M(p_{2})}\delta_{\rm in}(\boldsymbol{p}_{1})\delta_{\rm in}(\boldsymbol{p}_{2})\,.$ (3.8)
The contributions from higher-order PNG can be obtained through an analogous procedure. This expression allows for a straightforward incorporation of arbitrary-order PNG into the partition functions via appropriate corrections to the initial matter field $\delta_{\rm in}(\boldsymbol{k})$. For example, when we want to generalize the matter partition function presented in [23]
@@
where $\mathcal{P}[\delta_{\rm in}]$ is the initial probability distribution, which is usually assumed to be Gaussian, and $P_{\epsilon_{m}}(k)$ denotes the power spectrum of the noise in the matter density field, which is used to cancel the UV dependence of the loop diagrams [23]. At this point, all that is required is to substitute the non-Gaussian statistics of $\delta_{\rm in}$ eq. (3.8) into the probability distribution
@@
$-f_{\rm NL}\int_{\boldsymbol{p}_{1},\boldsymbol{p}_{2},\boldsymbol{p}_{3}}\left[\hat{\delta}_{D}(\boldsymbol{p}_{123})K_{\rm NL}(\boldsymbol{p}_{2},\boldsymbol{p}_{3})\frac{M(p_{1})}{M(p_{2})M(p_{3})}\frac{\delta_{\rm in}(\boldsymbol{p}_{1})}{P_{\rm in}(p_{1})}\delta_{\rm in}(\boldsymbol{p}_{2})\delta_{\rm in}(\boldsymbol{p}_{3})\right]\,.$ (3.10)
@@
$\times\exp\left\{-\frac{1}{3!}\,f_{\rm NL}\int_{\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2}}\hat{\delta}_{D}(\boldsymbol{k}+\boldsymbol{p}_{12})B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\delta_{\rm in}(\boldsymbol{k})\delta_{\rm in}(\boldsymbol{p}_{1})\delta_{\rm in}(\boldsymbol{p}_{2})\right\}\mathcal{P}[\delta_{\rm in}]\,$, (3.11)
@@
Note that the sign convention adopted here is consistent with [38], but opposite to that in [40]. [^7] In fact, both sign conventions are equally valid and yield the correct $n$-point statistics, as explicitly shown in eq. (4.2) of [38] and eq. (B.5) of [40]. The difference between them merely corresponds to a redefinition of the “coupling parameter” $B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\rightarrow-B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$. After matching with the $n$-point correlation functions, this distinction does not affect any physical predictions. In this context, the coupling parameter $B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$ does not exhibit symmetry under cyclic permutations of $k,p_{1},p_{2}$, but instead retains symmetry only under the exchange of $p_{1}$ and $p_{2}$. At the same time, $\mathcal{P[\delta_{\rm in}]}$ is still modeled as a Gaussian probability distribution, despite the fact that the underlying primordial statistics are no longer Gaussian. It is evident that, after incorporating PNG, the leading-order correction in $f_{\rm NL}$ effectively manifests as an additional interaction term in the partition function. Moreover, as illustrated in [40], this term can be absorbed into the term proportional to $J$ after performing a shift of the integration variable, which is consistent with the arguments in quantum field theory. Following the same procedure, we can derive the expression for the galaxy partition function with PNG
@@
$\times\exp\left\{\int_{\boldsymbol{k}}\left[\frac{1}{2}P_{\epsilon_{m}}(k)J(\boldsymbol{k})J(-\boldsymbol{k})+J_{g}(\boldsymbol{k})\delta_{\rm g,fwd}[\delta_{\rm in}](-\boldsymbol{k})+J(\boldsymbol{k})\delta_{\rm fwd}[\delta_{\rm in}](-\boldsymbol{k})+\cdots\right]\right\}$
$\times\exp\left\{-\frac{1}{3!}\,f_{\rm NL}\int_{\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2}}\hat{\delta}_{D}(\boldsymbol{k}+\boldsymbol{p}_{12})B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\delta_{\rm in}(\boldsymbol{k})\delta_{\rm in}(\boldsymbol{p}_{1})\delta_{\rm in}(\boldsymbol{p}_{2})\right\}\mathcal{P}[\delta_{\rm in}]\,.$ (3.13)
Compared to eq. (3.11), we have not only extended the current to separately correspond to the matter and galaxy fields, but have also generalized the stochastic term. This results in (a) the stochasticity for galaxies $P_{\epsilon_{g}}\sim k^{0}$, (b) the cross stochasticity between galaxies and matter $P_{\epsilon_{g}\epsilon_{m}}\sim k^{2}$, and (c) the matter stochasticity $P_{\epsilon_{m}}\sim k^{4}$ (where $\sim$ represents their leading-order contributions). When evaluating the joint likelihood in the large-scale limit, the matter stochasticity can usually be safely omitted, since it involves higher-order momentum dependence, which allows us to take $P_{\epsilon_{m}}\rightarrow 0$ [23].
In summary, in both cases, the effects of PNG can equivalently be reformulated as an interaction term under Gaussian initial conditions, thereby enabling the direct application of all analyses developed within the Gaussian framework. For example, in the current context, the matter likelihood can still be expressed as the functional Fourier transform of the matter partition function
@@
Here, we provide a brief comparison between the path-integral formulation and the standard perturbation theory (SPT) approach (see, for example, in [35]) in the context of incorporating PNG. In SPT, the introduction of PNG leaves the dark-matter equations of motion unchanged; instead, it modifies the initial conditions from which the perturbative expansion begins. At leading order, for example, one must specify a non-trivial initial bispectrum and include its contributions in the loop-level matter power spectra [38]. By contrast, in the path-integral formulation, although the underlying physics is the same, this framework admits a more economical, phenomenological treatment. As shown in eq. (3.10), PNG changes the initial probability distribution $\mathcal{P}[\delta_{\rm in}]$, but its effects can be absorbed into the partition functions as an effective interaction vertex (see eqs. (3.11) and (3.13)). The initial distribution can therefore still be treated as Gaussian, while all PNG effects are encoded in a single additional vertex, greatly simplifying the likelihood calculation. Moreover, the coefficients of these vertices are fixed by requiring that the path-integral formulation reproduces the same statistical properties obtained from the equation-of-motion approach. In the next section we will implement this strategy, use the saddle-point expansion to compute both likelihoods, and show how PNG introduces new, non-trivial contributions to the conditional likelihood.
@@ 3.3 Dimensional analysis
As we can see, the expressions for the two likelihoods, eqs. (3.14) and (3.15), involve numerous terms to be handled. Evaluating these likelihoods directly with all terms included poses a considerable challenge. Therefore, in this subsection, we perform a dimensional analysis to investigate the relative contributions of each term to the likelihoods at large scales. This can be achieved by performing a momentum rescaling analogous to the procedure outlined in [23], with $\boldsymbol{k}=b\boldsymbol{k}^{\prime}$. Then, in the limit $b\rightarrow 0^{+}$, the behavior of different terms under this rescaling will reveal their relevance (or irrelevance) to the large-scale universe. Under this rescaling, the integration measure ${\rm d}^{3}k$ changes to $b^{3}{\rm d}^{3}k^{\prime}$, and the power spectrum $P_{\rm in}$ scales as $P_{\rm in}(k)=b^{n_{\delta}}P_{\rm in}(k^{\prime})$, where $n_{\delta}$ is the order of momentum in the initial power spectrum. Then, the terms in the partition functions that are quadratic in $X_{g}$ and $\delta_{\rm in}$ remain invariant if we redefine
@@
These are the results of the rescaling in Fourier space, and the scaling behavior of these operators in real space can be obtained via the inverse Fourier transform
@@
To determine the rescaling of the non-Gaussian effective vertex, it is also necessary to study the scaling behavior of the non-Gaussian parameter $B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})$ in eq. (3.12). We begin in Fourier space, given the asymptotic behavior of the transfer function $T(k)\rightarrow 1$ in the limit $k\rightarrow 0$, it can be treated as dimensionless with respect to momentum at large scales and accordingly exhibits a trivial scaling behavior, namely $b^{0}$. In this way, $M(k)$ can be approximated to have a scaling behavior of $k^{2}$ in the relevant regime. Meanwhile, for the lowest-order momentum-dependent PNG, corresponding to the spin-0 case discussed in [13], the associated kernel function can also be considered momentum dimensionless. After substituting the momentum dependence of the initial power spectrum, the rescaling behavior of this parameter in real space is given by
@@
Based on these expressions, we can assess the contributions of the different terms in the partition function eq. (3.11) at large scales. We present the result of the rescaling of these terms and the evolution of their scaling dimensions with respect to $n_{\delta}$ in Table 1 and Figure 1, respectively. It is evident that for $n_{\delta}=-1.7$, both stochastic terms $P_{\epsilon_{g}}JJ$ and $P_{\epsilon_{g}\epsilon_{m}}J_{g}J$ exhibit rescaling with a higher-order dependence on $b$ relative to the PNG effective term, thereby yielding subdominant contributions at large scales. This is precisely why, in [40], the authors were able to disregard the stochastic terms and focus primarily on the non-Gaussian contributions when deriving the renormalization group equations incorporating PNG. At the same time, we can observe that in our case, the $P_{\epsilon_{g}\epsilon_{m}}J_{g}J$ term generally contributes the least to the partition function. This is because the leading-order momentum dependence of $P_{\epsilon_{g}\epsilon_{m}}(k)$ scales as $k^{2}$, resulting in a stronger suppression compared to other operators. Consequently, it is reasonable to neglect the higher-order contributions from $P_{\epsilon_{g}\epsilon_{m}}$ when calculating the likelihoods. The perturbative properties of $f_{\rm NL}$ and $P_{\epsilon_{g}\epsilon_{m}}$ will constitute the central framework for our detailed calculations in the next section, enabling us to extract the leading-order additional terms associated with the non-Gaussian contributions.
In fact, higher-order terms such as $JJJ$ should also be included in the construction of the partition functions to ensure the theory’s completeness under renormalization. However, in the present analysis, we neglect these terms for one primary reason: their contributions are negligible at large scales. To substantiate this claim, we refer to the dimensional analysis in [39], where the authors systematically investigate the dimensions of $J^{n}$ terms in real space. For a generic coupling between $J^{n}$ and an arbitrary operator $O[\delta]$, the effective vertex in the partition function and the dimension of its coefficient take the following form
$S_{\rm eff}\propto C_{O}^{(n)}(\Lambda)\int_{\boldsymbol{x}}[J(\boldsymbol{x})]^{n}O[\delta](\boldsymbol{x})\,,\qquad\qquad[C_{O}^{(n)}]=3-3n-d_{O}\,$, (3.20)
where $d_{O}$ is the dimension of the operator $O[\delta]$. As evident from this expression, the dimension of the coefficient decreases rapidly with increasing $J$-power. For instance, in the case of $JJJ$ term, this corresponds to $n=3$ and $d_{O}=0$, such that $C^{(3)}]=-6$. Consequently, the dimensionally consistent form is $\frac{C^{(3)}}{R^{6}}JJJ$, where $R$ represents the physical scale of interest (which is very large in our case). Therefore, higher-order terms are thus increasingly suppressed at large scales, justifying the neglect of terms beyond $J^{3}$ in likelihood computations.
As we have seen, two quantities in the partition functions eqs. (3.13) and (3.11) are suppressed at large scales: the cross stochasticity term $P_{\epsilon_{g}\epsilon_{m}}$, suppressed due to its higher-derivative nature, and the PNG parameter $f_{\rm NL}$, suppressed by the small amplitude of PNG. We now provide a brief estimate of their magnitudes. We first consider the $P_{\epsilon_{g}\epsilon_{m}}$ term, for which the scale dependence can be factorized following the arguments in [26] (this term corresponds to the $P_{\epsilon_{g}}^{\{2\}}$ term in that literature),
$P_{\epsilon_{g}\epsilon_{m}}(k)\sim AR_{*}^{2}k^{2}\,$, (3.21)
where $R_{*}$ is the typical nonlocal scale of galaxy formation, $A$ is some dimensionless constant and $k$ is the wavenumber corresponding to the scale in our study. In the context of our analysis, $R_{*}$ can be naturally identified with the nonlinear scale, $R_{*}\sim 1/k_{\rm NL}$. Substitute this into the expression for $P_{\epsilon_{g}\epsilon_{m}}$ in eq. (3.21), we can obtain
$P_{\epsilon_{g}\epsilon_{m}}(k)\sim A\left(\frac{k}{k_{\rm NL}}\right)^{2}\sim\left(\frac{10h^{-1}\text{Mpc}}{1000h^{-1}\text{Mpc}}\right)^{2}\sim 10^{-4}\,$, (3.22)
where we have fixed the nonlinear scale to $10h^{-1}\text{Mpc}$ and the scale of interest to $1000h^{-1}\text{Mpc}$. It should be note that the magnitude of $P_{\epsilon_{g}\epsilon_{m}}$ is scale-dependent. On the other hand, the treatment of the non-Gaussian contributions is a little more subtle. Taking the local-type non-Gaussianity model as an example, based on the most recent constraints from the Planck satellite, $f_{\rm NL}^{\rm local}=-0.9\pm 5.1,$ [59], the parameter $f_{\rm NL}$ may be consistently taken to be of order $\mathcal{O}(1)$. Then this analysis indicates that the large-scale contribution of $P_{\epsilon_{g}\epsilon_{m}}$ is subdominant compared to that from $f_{\rm NL}$. It is therefore reasonable to neglect the higher-order terms in $P_{\epsilon_{g}\epsilon_{m}}$, as was done in [23]. However, current observations do not entirely exclude the possibility of a small but non-vanishing level of non-Gaussianity, as generally predicted by slow-roll inflationary scenarios [5], e.g. $f_{\rm NL}\sim\mathcal{O}(10^{-2})$. Therefore, We do not rule out the possibility that the contribution of $f_{\rm NL}$ may be subdominant compared to that of the derivative term (note that the magnitude of $P_{\epsilon_{g}\epsilon_{m}}$ can be also be altered by adjusting the scale under consideration). In line with the majority of existing literature [13, 21, 40], we choose to first focus on the terms proportional to $f_{\rm NL}$, as they are the most relevant for probing PNG. At the same time, we treat the $f_{\rm NL}P_{\epsilon_{g}\epsilon_{m}}$-order terms as the corrections to the leading-order PNG contribution and neglect all contributions of order $f_{\rm NL}^{2}$ and higher. As will be demonstrated in the following section, relative to the case of Gaussian initial conditions, the first-order saddle-point corrections to the conditional likelihood is proportional to $f_{\rm NL}P_{\epsilon_{g}\epsilon_{m}}$, whereas the corrections due to saddle-point fluctuations scales as $f_{\rm NL}$.
4 Precision calculation of the EFT likelihood